OpenStudy (anonymous):

I need help with the following sum: sum (y^2(lambda^y*e^-lambda/y!)) from 0 to infinity I know the answer is lambda+lambda^2. But I need to know why.

OpenStudy (anonymous):

What's the index on the sum? y?

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

Wait, is it: \[ \sum_{y=0}^\infty y^2\frac{\lambda^y\cdot e^{-\lambda}}{y!} \]

OpenStudy (anonymous):

@wio yes. It is.

OpenStudy (anonymous):

Hmm, well I know: \[ e^x=\sum_{k=0}^{\infty}\frac{x^k}{k!} \]

OpenStudy (blockcolder):

Start with \[e^\lambda=\sum_{y=0}^\infty \frac{\lambda^y}{y!}\] Differentiate this wrt lambda to get \[\lambda e^{\lambda-1}=\sum_{y=0}^\infty y\frac{\lambda^{y-1}}{y!}\] Multiply both sides by lambda to get \[\lambda^2 e^{\lambda-1}=\sum_{y=0}^\infty y\frac{\lambda^y}{y!}\]

OpenStudy (blockcolder):

Do this once more to get a closed form for \[\sum_{y=0}^\infty y^2 \frac{\lambda^y}{y!}\] then multiple both sides by e^-lambda.

OpenStudy (anonymous):

\[ \frac{d}{d\lambda }\lambda ^y = \lambda ^y\ln(\lambda) \]

OpenStudy (blockcolder):

Correction: In the second line of my first post, that should be \[e^\lambda=...\]

OpenStudy (blockcolder):

No, it's \[\frac{d}{dy}\lambda^y=\lambda^y \ln{\lambda}\]

OpenStudy (anonymous):

Oh, I see.

OpenStudy (anonymous):

Okay?

OpenStudy (blockcolder):

In the end, you get \[\lambda(\lambda+1)e^\lambda e^{-\lambda}=\sum_{y=0}^\infty y^2 \frac{\lambda^y}{y!}e^{-\lambda}\]

OpenStudy (blockcolder):

The LHS can be simplified into \(\lambda(\lambda+1)=\lambda^2+ \lambda\).

OpenStudy (anonymous):

Thank you!